You are here: Foswiki>PartnerObserving Web>TechNightLogs>2009March21UT (21 Jul 2009, NormCushing)Edit Attach

21 March 2009 UT - LBC Technical PartnerObserving

Observers: O. Kuhn (@LBT), J. Hill (Tucson)
Telescope Operator: D. Gonzalez Huerta

Problems with the NFS disk at the beginning of the night.

In the afternoon I (Olga) analyzed the red focus sequences to determine filter focus offsets. There was scatter, even though the seeing was excellent and stable, and the fits were good (some Blue fits were not so good, though.) Two to four sequences per filter were obtained and collimation was done frequently, every 2-3 sequences or every half hour. A list of the relevant images and the filters and OB names, is at ListofFocusData and the results are at FocusResults. The results were entered in the configuration file and a new line was added to the Focus Offsets between Red Filters table in the Engineering Section of the LBC User Manual.

Data were obtained also to produce new collimation lookup tables for Blue and Red. The collimated pupil images obtained before running focus sequences were used in addition to several pupil images obtained during the last hour of the night. The new collimation lookup tables were copied to /home/telescope/TCS/Configuration/PSF/new. All global offsets are zero. Last night we did not experience problems with running into any of the X, Y, or Z limits. Temperature was about 4C.

Following are links to the collimation model plots for SX and DX Model Plots.

~21:00 LBC UI in a comatose state - It ran up without problems, and allowed me to upload and play an OB. But nothing happened - I loaded a 'fast OB' and clicked play, button greyed out, but the telescope did not slew and there was nothing updating on the user interface. The camera status on the OB execution page showed disabled, but on power control all ready. lbckill/lbcstart allowed me to restart. I did turn it on hours ago (while it was being filled, in case that might have caused a problem); also I did not do anything with it until about 21:00 when we opened.

21:10 Slewing to 77deg elevation to collimate. dofpia,/first The /first will send a larger initial Z11 (red 1000, blue 1200).

21:20 No data appearing in Repository. LBC archive machine affected by NFS. John rebooting lbcarchive2 remount disks.

21:30 dofpia,/backout since dofpia had crashed because it could not find data files. Should have run with with the /first keyword, but instead ran it two times to remove most of corrections.

21:37 Data still not appearing in Repository, though we do see them in lbcarchive (in lbtsdt directory- note the OBSERVER keyword should be changed to Calibration in the version of the OB file called by dofpia). Tracked down to the fact that a lbcarchive2 cannot scp files to Repository. Issuetrak #2081

22:33 Issue #2081 solved. dofpia running. Taking a while to converge - red pupils in particular look 'blobby' - DX glass and ambient temps differ by about 1.5-1.1C.

23:00 MV had tripped off - fixed. Going to a pointing star to adjust pointing. ACT0290 at elevation 80 deg. Collimate there. The Red pupils consistently show a ring. David is giving -200nm Z22 to DX. Last pupil image before giving Z22 is lbcr.20090321.061044.fits - to add to manual if indeed it is an example of positive Z22.

23:14 restart dofpia.

note output from lbcfpiaBlueLog shows that while Blue converged (054249), it took and analyzed extra files while Red was still converging. The Blue convergence did not hold. This gives an idea of the scatter that can be expected in the coefficients. Assuming the ratio: 1000nm Z4 for every 0.0379mm Z, 400nm corresponds to about 15 microns Z.
  -27    -95   -192     -2      3   -261    y  1.44  0.19                   /Repository/lbcb.20090321.054249.fits  V-BESSEL
   452   -282    133     84     39    -77    y  1.49  0.22                   /Repository/lbcb.20090321.054518.fits  V-BESSEL
   314    401    -56    -32    -41    114    y  1.54  0.20                   /Repository/lbcb.20090321.054717.fits  V-BESSEL
  -409   -107     55     24     27    -23    y  1.38  0.18                   /Repository/lbcb.20090321.054922.fits  V-BESSEL

23:30 Red side was not converging well. Stopped and slewed to a new pointing star at higher elevation, ACT0322 (elev=83.4 deg). Exited/restarted IDL to get new dofpia version.

23:45 ACT0322 r is 065232, but it took a very long time to start the red camera exposure. CA and IE are: CA = -28.95 and IE= -23.17. New values: NEW CA, IE: -17.2 -40.6. Check with 065526. For this image, ptadjust gives:

 Star at pixel coordinates: 1069.527, 2881.345   (rotator center is at 1080, 2891)

  R.A. Dec.Offsets: 2.345952, -2.16272

  NEW CA, IE: -19.5 -42.8
OK

23:57 Check co-pointing at very high elevation. Going to WT10_337.

00:02 New DOFPIA ready. If you type "S", it will stop the iteration and give the 'exit' corrections to Z11 and Z4. +mirradjust+ gives the result for SX global offsets needed to copoint at this star, elev = 82
Need to adjust the SX mirror by 
dX = 0.99 mm dY = 1.12 mm dRX = 23.47 arcsec dRY = -20.79 arcsec 

00:25 New CA = -27.0 and NEW IE=-35.0 entered based on 072254. Now take another exposure to correct co-pointing. Elev=79.9 deg.
Need to adjust the SX mirror by 
dX = 1.23 mm dY = 1.18 mm dRX = 24.78 arcsec dRY = -25.77 arcsec 
David entered these values. Ignore next set, taken while mirror moving. Check 072945,072946.
delta_az = 12.39 arcsec delta_el = -13.23 arcsec 
Need to adjust the SX mirror by 
dX = 0.60 mm dY = 0.64 mm dRX = 13.49 arcsec dRY = -12.63 arcsec 
Did it move so much already? Am taking another of this star 073802,073800 = b,r, before going to higher elevation.
delta_az = 2.13 arcsec delta_el = -3.88 arcsec 
Need to adjust the SX mirror by 
dX = 0.10 mm dY = 0.19 mm dRX = 3.96 arcsec dRY = -2.17 arcsec 
No but there is some scatter in my measurements of the star or in positions with changing collimation.

00:41 Go to ACT0324. 074348, 074349 are taken before collimation. Collimate. Lost hole on red pupils - David going to remove the Z22, so he's giving Z22 = +200nm to DX. Collimate and take data at ACT0324, elev=87 deg. mirradjust yields:
dX = -0.51 mm dY = -0.07 mm dRX = -1.50 arcsec dRY = 10.79 arcsec 

Combining with previous set we get a new set which David is applying: X=0.72 Y=1.11 RX=23.28 and RY=-14.98

Take one more pair. Elev=85. These 080531 b,r give:
Need to adjust the SX mirror by 
dX = 0.19 mm dY = 0.09 mm dRX = 1.86 arcsec dRY = -4.04 arcsec 
Good. So above set of GO's will co-point at elev = 85. I am putting them into SXPMLBCCollimation.dat on TCS/Configuration (we did this and re-initialized the lookup table). We are collimating again at the same position before taking another pair of images to check, but red is not converging and I think co-pointing is falling off as we move to lower elevation (now at elev =81). Blue had converged quickly, though it did get a spike in Z4.
  450      0      0      0      0    500    y  0.00  0.00                     dofpia_sph3_adjustment
 -2399    -73   -142    115    215   -411    y  0.52  0.99                   /Repository/lbcr.20090321.081611.fits  r-SLOAN
   669    153    -55   -164     15     99    y  0.55  0.30                   /Repository/lbcr.20090321.081805.fits  r-SLOAN
  -609    -74    219     47    -44   -326    y  0.55  0.33                   /Repository/lbcr.20090321.081952.fits  r-SLOAN
   959     53   -116    -56    -93    251    y  0.59  0.42                   /Repository/lbcr.20090321.082141.fits  r-SLOAN
  -482   -118     69    244    195    213    y  0.60  0.29                   /Repository/lbcr.20090321.082327.fits  r-SLOAN
  -932    -11   -171   -352   -226   -333    y  0.50  0.47                   /Repository/lbcr.20090321.082515.fits  r-SLOAN (after this I gave "S" to stop and send final correction)
  -225      0      0      0      0   -250    y  0.00  0.00                     dofpia_sph3_adjustment
  
 540      0      0      0      0    600    y  0.00  0.00                     dofpia_sph3_adjustment
 -2889    239    146   -146    -31   -160    y  1.02  1.15                   /Repository/lbcb.20090321.081612.fits  V-BESSEL
    59     12    -99    -12     56    -12    y  1.01  0.09                   /Repository/lbcb.20090321.081806.fits  V-BESSEL
    15     99    137    -25    -16    182    y  0.91  0.15                   /Repository/lbcb.20090321.081953.fits  V-BESSEL
  -571   -219    -44     52    -53    -49    y  1.02  0.25                   /Repository/lbcb.20090321.082142.fits  V-BESSEL
  -135    194    -76   -114     -5    -44    y  1.08  0.13                   /Repository/lbcb.20090321.082328.fits  V-BESSEL
    63      4     53     -5     40     21    y  0.93  0.09                   /Repository/lbcb.20090321.082516.fits  V-BESSEL
  -562      0      0      0      0   -625    y  0.00  0.00                     dofpia_sph3_adjustment

01:40 Slewed to lower elevation, 37 deg, to check X,Y limits of SX,DX. Looks like the Red pupils almost converged on first image, the Blue has a large focus offset (4740), but Z5-Z11 not too large. 01:58 Converged at elev = 36.6
085658 (R), 085659 (B) - not close to any mirror limits, temp is about 3C.
pair of images in-focus 085929 (R), 085929 (B)

02:00 Slew to Virgo cluster for multi-filter OB to test guiding/dithering. /home/LBTO/0ObservingBlocks/virgocluster.ob I stopped the OB, because I forgot to ask David to slew and the OB would have taken. Red converged in 2x and Blue in 2x! Elev = 63 deg
090747(R), 090748(B)

02:10 Virgo observation. Blue has guide star but Red (nb filter) does not. raises issue #1989 again I should have pre-selected dithers to put star on red guide chip.

Red images at first dither:
  • There is a large gradient in IQ from good at top of chip 2 (~3.5 pix fwhm) to bad (5.2 pix fwhm) at bottom of chip 2 in lbcr.20090321.091336.fits Y-FAN. (guide star faint).
  • Similar gradient for lbcr.20090321.091102.fits F972N20 (not guided)

Blue images at first dither:
  • lbcb.20090321.091104 has fwhm about 5.1 up/down along chip 2 SDT-Uspec
  • lbcb.20090321.091341 also has fwhm about 5.1 up/down on chip 2 g-SLOAN

So far we see image gradient in Red, not in Blue - will try other filters. There is also this uneven scattered light gradient in both Red filters (F972 and Y).

02:40 dofpia again. Elev = 57 deg. Red converged on first iteration, Blue took 3x.
094709(R), 094710(B) 
02:50 Virgo cluster again at B,r and V,i. FPIA predicted seeing 1.3(B) and 0.7(R) but guide stars have 2.6"(B) and 1.9" (R)
  • V on R 094952: slight gradient top/bottom but not so much as before. fwhm 3.7 at top of chip 2 to 4.6 at bottom. - later image 101210 has fwhm 3.5-4 and better at bottom instead of top.
  • i-S on R 095227 gradient 3.3 fwhm at top and 4.4 fwhm at bottom. 101504 has fwhm ~3.5
  • B on B 094954 fwhm 4.6-4.7 --- later image 101210 has fwhm ~4.3 so collimation holding ok for 22 min at least
  • r-S on B 095229 fwhm 4.1 up/down along chip 2, 101516 has fwhm ~ 4.2 but looks like at the bottom of chip 2 stars there is some coma.

03:21 Collimate again at position. Blue converged in 1 iteration, Red in 2. Elev = 54 deg
102506(R), 102507(B)
Virgo at U,V (B) and z,I,R (R)

  • 102810 z-SLOAN images: look elongated: fwhm 3.3 at top and elong~0.08 but 4.3 at bottom, elong~0.12
  • 103049 I-BESSEL images: 3.6 at top and pretty round, but 4.1 at bottom and elongated in same sense as for 102810
  • 103211 R-BESSEL images: ~4.3 fwhm

  • 103057 V-BESSEL: 3.6-4 some elongation
  • 103422 U-BESSEL: 4.7

03:45 DX Panicked At 03:44 one of the actuators became disabled - panic occurred during 104351.

03:54 DX back up. Collimating. Blue converged in 2x, Red didn't and there was a delay in images getting to Repository, then next an error which may also result from slow data transfer:
Waiting for OB to finish and data files to appear in /Repository
Processing: /Repository/lbcr.20090321.110550.fits
FITS_READ ERROR: READU: End of file encountered. Unit: 101
         File: /Repository/lbcr.20090321.110550.fits
When collimating, I noticed error messages, "Error on TCSInitialize()", e.g.
Sending last Sph3 (Z11) correction for RED channel      -322.644

-------------------------
 TCS Send Wavefront program
-------------------------

Error on TCSInitialize()
Exit 1

DOFPIA continued, and appeared to be sending corrections, 2 more lines in both blue and red, but David told me that the last corrections sent to the mirror were Z11=-22 for Blue and Z11=-143 for Red, which correspond to image numbers 110939 (B) and 110925 (R). When the Error appears, the corrections are not sent even though DOFPIA continues to take and analyze pupil images. Issuetrak to be filed.

   540      0      0      0      0    600    y  0.00  0.00                     dofpia_sph3_adjustment
 -1632    105   -406   -461     17    -22    y  1.23  0.69                   /Repository/lbcb.20090321.110939.fits  V-BESSEL <--- last correction sent to SX
  -293   -216    -51    -30     54   -262    y  1.38  0.23                   /Repository/lbcb.20090321.111357.fits  V-BESSEL
  -705      0      0      0      0   -783    y  0.00  0.00                     dofpia_sph3_adjustment

   450      0      0      0      0    500    y  0.00  0.00                     dofpia_sph3_adjustment
  -932     17   -191    123    479   -143    y  0.80  0.45                   /Repository/lbcr.20090321.110925.fits  r-SLOAN <--- last correction sent to DX
   184    -72    163    300   -383    -92    y  0.71  0.26                   /Repository/lbcr.20090321.111356.fits  r-SLOAN  
  -290      0      0      0      0   -322    y  0.00  0.00                     dofpia_sph3_adjustment

I restarted DOFPIA and it is sending corrections again. A network problem? Just before I had had the slow image transfer.The error in sending corrections happened earlier tonight or yesterday and also cleared up by itself.

04:25 recollimate - at elevation 34 deg, but we are not close to mirror limits. It is taking a while to collimate, so I am stopping it with "S". Red had converged but not blue. Taking a few images of Virgo from the OB when the mirror panicked.
 Red collimated pupil image is 113315 

04:40 Loaded OB and clicked Play, but camera spontaneously disabled. I admit, I have stopped an OB or two during the night... It is automatically restarting. I tried to load the same OB, virgocluster3.ob, but again camera disabled, just when I was loading it. Restarting automatically. The problem was that I had edited the dither sequences in the OB and not updated the number of dithers to be consistent. I did this and it is loading and running now. Paused OB at 04:55. Finishing 115521 and 115525. Restart OB (* check - where does it restart as far as dither pattern?* 0,-20,40 and 0,40,-20).

05:15 Slew to elev=76deg and collimate. zBsuperfoc sequence and then Ir superfoc, but on this one we have a warning on Red. *Filters hardware failure. filter is more than 10 times the maximum allowed distance from it nominal position [src/filters/filters.c:2153]FILTERS WHEEL#1*

05:36 Trying another OB, rUs superfoc. Filter is working now.

05:41 Preparing for flats. Changed typical exposure time to 1sec and minimum scientific exposure time to 0.5 so that tracker images will be obtained during our twilight flats. These are to address Stefano's request in issuetrak #1393.Turned camera off/on, but Blue trackers did not come up. Turning off/on again, OK now.

05:58 Plan to try V,Y flats (blue V was saturated, Y 8000 cnts) and then U-B, F972N20. It was too bright for V, so I tried F972N20 and U. OK for these, but still with the edits I made to the conf files and with exposure times of1sec for the science chip images and 1 sec (typical) for trackers, we are not getting tec chip images. I forgot to decrease the tracker delay time (TrackerLoopStartDelay is currently 10sec). Disabled B since U counts were high, to get some F972N20 flats at least.

06:08 Closing. I am restoring the conf files.

-- JohnHill - 21 Mar 2009
Topic revision: r4 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback